When CO2(g)CO2(g) is put in a sealed container at 730 KK and a pressure of 10.0 atm and is heated to 1420 KK , the pressure rises to 24.1 atm. Some of the CO2CO2 decomposes to COCO and O2O2.


Calculate the mole percent of  CO2 that decomposes.


Express your answer using two significant figures.


%CO2 =




1
Expert's answer
2021-06-03T05:27:03-0400

Ideal Gas Law:

pV=nRT

R = 0.08206 L×atm/mol×K

"10.0 \times V = n \times 0.08206 \times 730 \\

10.0 \times V = 59.90n \\

n = 0.1669V"

After decomposition:

"24.1 \times V = n^* \times 0.08206 \times 1420 \\

24.1 \times V = 116.52 \times n^* \\

n^* = 0.2068V"

(which is the total amount of moles of the gases)

Let's suppose, V = 1L, so:

"n=0.1669 \\

n^*=0.2068"

For the decomposition reaction, we can do a reaction table:

2CO2 → 2CO + O2

0.1669____ 0___ 0

-2x______ +2x__ +x

0.1669-2x __2x__x

"n^* = 0.1669 -2x + 2x +x \\

0.1669 + x = 0.2068 \\

x = 0.0399"

So, "2 \times 0.0399 = 0.0798" mol of CO2 decomposes, the percent is:

"\frac{0.0798}{0.1669} \times 100 \;\% = 47.81 \; \%"

Answer: 47.81 %